I need help to solve this "Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?"
Answer by checkley77(12844)

SHORT=LONGEST/2+30
MIDDLE=LONGEST-2
LONGEST+(LONGEST/2+30)+LONGEST-2=378
L+L/2+30=L-2=378
2.5L=378-30+2
2.5L=350
L=350/2.5
L=140 YDS. IS THE LONGEST SIDE.
S=140/2+30=70+30=100 YDS. IS THE SHORTEST SIDE.
M=140-2=138 YDS. IS THE MIDDLE SIDE.
PROOF:
140+100+138+378
378=378